how do you simplfy i to the 91th power
Answer by stanbon(48538)

how do you simplfy i to the 91th power
---------
i^4 =1 and 91 = 4(22)+3
So i^91 = i^(4*22+ 3) = i^(4*22)*i^3 = 1^22*i^3 = i^3 = -i
===========
Cheers,
Stan H.